At what temperature will the molecule of an ideal gas have twice the the rms speed they have at 20°c

1
Expert's answer
2021-11-03T10:19:58-0400

The speed is proportional to the square root of temperature:


"v_\text{rms}=\sqrt\frac{3RT}{M},\\\space\\
u=2v_\text{rms}=2\sqrt\frac{3RT}{M}=\sqrt\frac{3R(4T)}{M},"

so, the temperature is 4T=4(273+20)=1172 K or 899°C.